OpenStudy (lgbasallote):

i need to learn derivatives of transform...can someone give me a demo? for example \[\text{Find} \; \mathcal L \lbrace t^2 \sin kt \rbrace\]

OpenStudy (anonymous):

I'm learning from this question :)

OpenStudy (lgbasallote):

good for you! :D

OpenStudy (anonymous):

waiting for the answer .. sami , please :)

OpenStudy (lgbasallote):

btw...please dont post a full solution...i learn better when it's step by step :DD

OpenStudy (anonymous):

yes let me try. \[\Large L[t^nf(t)]=(-1)^n \frac{d^n}{ds^n}[F(s)]\] ok

OpenStudy (lgbasallote):

what does that mean?

OpenStudy (anonymous):

this means first find the Laplace of f(t) here f(t)=sin(kt) what is its Laplace.?

OpenStudy (anonymous):

L(sin kt)=k/(s^2+k^2) \[L(t \sin kt)=--d/ds(k/s ^{2}+k ^{2})\]

OpenStudy (lgbasallote):

uhh let's do this slowly... what does this mean \[\mathcal L \lbrace t^n f(t) \rbrace\]

OpenStudy (lgbasallote):

t^n is t^2 and f(t) is sinkt?

OpenStudy (anonymous):

\[L( t sinkt)=-d/ds(k/s ^{2}+k ^{2})\]

OpenStudy (anonymous):

its means Laplace when a function f(t) is multiplied by t^n here n=2 in your post . so first you need Laplace of f(t)=sin(kt)

OpenStudy (lgbasallote):

...so im right?

OpenStudy (anonymous):

yes you are right

OpenStudy (lgbasallote):

nice...

OpenStudy (lgbasallote):

so \[\large (-1)^n \frac{d^n}{ds^n} [F(s)]\] is this a formula?

OpenStudy (anonymous):

yes it is formula . you need first of all the Laplace of sin(kt ) once you find the Laplace of sin(kt), then take its nth derivative with respect to s (here n=2 so second derivative) then multiply the result of derivative with (-1)^n, in this case (-1)^2 =1

OpenStudy (lgbasallote):

that is the EXACT formula?

OpenStudy (anonymous):

yes it is !

OpenStudy (lgbasallote):

my notes says something like \[f'(s) = \int _0^\infty (-t) e^{-st} f(t) dt\] are those the same? btw...im not too sure about this formula...my notes are a little vague

OpenStudy (anonymous):

your notes are talking about transform of derivative .it is correct. while the question you have posted is different .it needs the formula that i have provided.

OpenStudy (lgbasallote):

uhmm okay... F(s) = L{f(t)} corret?

OpenStudy (lgbasallote):

correct*

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

uhmm wait...when do i use the formula you gave and when do i use the formula in my notes?

OpenStudy (anonymous):

the formula i gave you always use it when you see a function is multiplied by t,t^2,t^3.... while the formula in your notes apply that whenever you encounter derivative . f',f'',f''' and you need their Laplace.

OpenStudy (anonymous):

@lgbasallote i just checked the formula you have mentioned is used during the proof of the formula i have provided !!!!!

OpenStudy (anonymous):

but you would not use that one. its end result is what i have mentioned. stick to that one.

OpenStudy (lgbasallote):

uhh so in layman's terms...?

OpenStudy (lgbasallote):

the formula you gave is applicable when i have L{t^n f(t)}?

OpenStudy (anonymous):

always.

OpenStudy (lgbasallote):

what do you mean always?

OpenStudy (anonymous):

always applicable when you have L{t^n f(t)}

OpenStudy (lgbasallote):

oh. it's only applicable when it's like that?

OpenStudy (anonymous):

*making notes*

OpenStudy (lgbasallote):

or it applies in some other specific forms too?

OpenStudy (anonymous):

yes it is applied to that form .and it has applications when you will need inverse Laplace .

OpenStudy (lgbasallote):

in layman's terms...?

OpenStudy (anonymous):

in Layman's terms whenever you encounter multiplication of any standard function by t and powers of t this formula will be the only option there to save you. for example L(te^t) L(t^3sin(t) L(t^4e^at) L(t^2cosh(t)) L(t^3e^(at)sin(t)) the only formula is the above one this is the only only and only way you will be able to get their Laplace.

OpenStudy (lgbasallote):

yes...but as far as getting laplace is concerned L{t^n f(t)} is the only time the formula you gave is applicable?

OpenStudy (anonymous):

for now yes only when L{t^n f(t)}. But you will see this as you go through your course.

OpenStudy (lgbasallote):

well anyway....on to the solution

OpenStudy (anonymous):

ok find Laplace of sin(kt) first. i hope you can do this ?

OpenStudy (lgbasallote):

n = 2 and f(t) = sin kt

OpenStudy (anonymous):

yes correct.

OpenStudy (lgbasallote):

\[F(s) = \mathcal L \lbrace \sin kt \rbrace \implies \frac{k}{s^2 + k^2}\]

OpenStudy (anonymous):

yes it is correct now differentiate this two times (because n=2) with respect to s.

OpenStudy (lgbasallote):

so \[\mathcal L \lbrace t^2 \sin kt \rbrace = (-1)^2 \frac{d^2}{ds^2} (\frac{k}{s^2 + k^2})\] ??

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

is the first derivative \[\frac{-k(2s+k^2)}{(s^2+k^2)^2}\]

OpenStudy (anonymous):

use Quotient rule . in the numerator it should be just -2s because k is constant.

OpenStudy (lgbasallote):

oh lol yeah..

OpenStudy (lgbasallote):

wait...shouldnt it be -2sk?

OpenStudy (anonymous):

yes it should be .

OpenStudy (lgbasallote):

is the second derivative \[\frac{-2k(s^2 + k^2)^2 + 8s^2k (s^2 + k^2)}{(s^2 + k^2)^4}\]??

OpenStudy (anonymous):

ok wait let me calculate as well !

OpenStudy (lgbasallote):

okay...

OpenStudy (anonymous):

i think you made a mistake i am getting \[\Large \frac{2k(3s^2-k^2)}{(s^2+k^2)^3}\] check it again.

OpenStudy (lgbasallote):

but how :O

OpenStudy (lgbasallote):

now i have \[\huge \frac{-2k (-3s^2 + k^2)}{(s^2 + k^2)^3}\] i think they're same right?

OpenStudy (lgbasallote):

seems you just simplified farther than i did

OpenStudy (anonymous):

ok after first derivative we have \[\Large \frac{-2ks}{(s^2+k^2)^2}\] i am always comfortable with product rule :P so we can write it as \[\Large -2ks(s^2+k^2)^{-2}\] now by product rule \[ \frac{d}{ds}(-2ks(s^2+k^2)^{-2})=-2k((s^2+k^2)^{-2}-2ks(-2)(s^2+k^2)^{-3}*2s\] can you simply it now

OpenStudy (lgbasallote):

so this is the answer for \[\huge \mathcal L \lbrace t^2 \sin kt \rbrace\] since (-1)^2 = 1

OpenStudy (lgbasallote):

i got it already lol

OpenStudy (anonymous):

ok

OpenStudy (lgbasallote):

it was just a difference in simplification

OpenStudy (lgbasallote):

so is this the final answer?

OpenStudy (anonymous):

yes it is the answer because multiplication by (-1)^2=1 will not make any difference it is final answer !!!!!!!! after 1 and half hrs :P

OpenStudy (lgbasallote):

wow cool..my book agrees :D

OpenStudy (lgbasallote):

and s>0 right?

OpenStudy (anonymous):

yes